OpenStudy (anonymous):

HELP HAHD......... (x+1)(X-1)=?????

OpenStudy (anonymous):

there are no solutions

OpenStudy (akshay_budhkar):

lol! polpak will suspend you for spamming hasir!!

OpenStudy (anonymous):

the answer is x^2 - 1^2 that is x^2 - 1

OpenStudy (anonymous):

@kushkawa explain plz

OpenStudy (anonymous):

maybe you should get out more.

OpenStudy (anonymous):

@satellite i am really stuck in this question

OpenStudy (anonymous):

ok , as we know that (a+b)(a-b) = (a^2 - b^2 ) hence here ( x+1)(x-1) = (x^2 - 1^2) = x^2 - 1

OpenStudy (akshay_budhkar):

stop joking hashir

OpenStudy (anonymous):

LOOOOOOOOOOOOOOOOOOOOOOOOOOL

OpenStudy (anonymous):

i m not jokin

OpenStudy (anonymous):

explain plzzzzzzzzzzzzzzzzzzzz

OpenStudy (saifoo.khan):

want help?

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

look the method is known as FOIL

OpenStudy (anonymous):

i wanna foil his plans

OpenStudy (saifoo.khan):

Can you please stop spammming, i want to explain!

OpenStudy (anonymous):

what is foil

OpenStudy (saifoo.khan):

FOIL is not the foil u have in @ your house

OpenStudy (anonymous):

tin foil?

OpenStudy (saifoo.khan):

its FOIL

OpenStudy (saifoo.khan):

http://en.wikipedia.org/wiki/FOIL_method

OpenStudy (anonymous):

Force of OIL

OpenStudy (saifoo.khan):

Right^

OpenStudy (akshay_budhkar):

guys stop it hashir what is your main aim!!!!!lol

OpenStudy (saifoo.khan):

his aim is the understand FOIL

OpenStudy (anonymous):

hashir bhaiya , have u got it

OpenStudy (saifoo.khan):

this is conspiracy^

OpenStudy (anonymous):

kya hashir bhaiya reply to do ,

OpenStudy (anonymous):

saifoo clic on hashir's name and look at previous questions asked

OpenStudy (saifoo.khan):

that is easy

OpenStudy (saifoo.khan):

this is srsly difficukt

OpenStudy (saifoo.khan):

difficult*

OpenStudy (anonymous):

Here we go again...

OpenStudy (anonymous):

well the question was too much hard right bhaiya , hahaha